Reports 1-1 of 1 Clear search Modify search
AdV-PAY (NE and WE Payloads)
mwas - 19:40 Wednesday 07 May 2025 (66736) Print this report
High 50Hz for WE DR coil

To test the idea that a small fraction of the ring heater current can be flowing through one of the mirror coils I have looked at the trend of these coil monitors before and after the mirror change, and before and after the ring heater was turned on, but without the interferometer locked. So that the monitors of the coil voltage are dominated by the mirror control corrections themselves.

Figure 1 shows that there is a change in the average voltage during the several days when the payload is not connected between Apr 7 and Apr 11. All of the voltage go back to the same level as before the payload is extracted, with the exception of the DR coil which stays at the same level. The time when the ring heater is turned on doesn't change the voltage seen by these coil monitors.

Figure 2 shows the spectra, both with the ring heater turned off, before (purple) and after (blue) the change. The DR coil monitor started to see a large 50Hz and harmonics signal on them after the WE change, while the other coils have remained the same. This suggest that there is cabling issue with the DR coil.  It would be good to find a more direct way of electrically check that coil, to see if one of the two cables or the shield is connected with something it shouldn't be.

Note that the other large lines around 350Hz, 1kHz and 2kHz are calibration lines applied to the coils: https://wiki.virgo-gw.eu/Commissioning/Calibration/HistoryOfCalibrationLineFrequencies.

 

Images attached to this report
Comments to this report:
mwas - 11:31 Monday 12 May 2025 (66752) Print this report

Figure 1. The increase in 50Hz harmonics on the WE DR coil started on Saturday April 12 in the morning around 07:16 UTC (9.16am LT).

Which after the final electrical checks: https://logbook.virgo-gw.eu/virgo/?r=66564

And at the beginning of the work of payload integration that happened on April 12, https://logbook.virgo-gw.eu/virgo/?r=66569

It could be that something got damaged during the integration, or could be a simple as the problem appeared only once the cables got connected to the right places. It would be nice to check the times when various tests were done last week, to see if the problem temporarily disappears during some of the test, when some of the cables are not connected.

Figure 2 During the check of injecting a square wave into the RH the 50Hz harmonics in the the DR coil have disappeared https://logbook.virgo-gw.eu/virgo/?r=66745

Figure 3 shows that there have actually been several changes in the past few weeks. In purple now after the ring heater was turned on, the problem is no longer present. In blue is the night before injecting a square wave into the RH, there is 50Hz and an extra low frequency noise, but no 150Hz (the extra 50Hz and noise are actualy present also on ather coils. And in red is the initial problems seen after installing WE in april.

 

Images attached to this comment
dattilo - 14:22 Monday 12 May 2025 (66754) Print this report
The "extra noise" noted in figures 2 and 3 is actually not noise, but rather the correction signal of the mirror's z-damper, wich uses the mirror coil.
In the attached plot you can see the relevant signals for this z-damper and the times of when it was on. We turned it off at the beginning of our test (https://logbook.virgo-gw.eu/virgo/?r=66745), so as not to mask the signal we wanted to measure.
Images attached to this comment
mwas - 15:35 Monday 12 May 2025 (66756) Print this report

Figure 1. Looking back in the data, the 50Hz harmonics I have seen in the DR coil after the WE replacement have disappeared on Apr 22, on the first day of the recovery a few hours after the gate valve was open. One the damper correction are started, a minute later the 150Hz, and higher odd harmonics of the 50Hz disappear. The 50Hz is masked by the high 50Hz correction of the damper correction.

Hence the issue with the 50Hz harmonics on the DR coil was never present once we had the interferometer, and this was a false leed for the questions of high magnetic coupling and 3Hz comb.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×